OpenStudy (brooke2353):

5+3+2 = 151022 9+2+4 = 183652 8+6+3 = 482466 5+4+5 = 202541 THEN ; 7+2+5 = ? lol will medal and fan

OpenStudy (anonymous):

do you play the blockheads brooke? i know someone with your name who dose

OpenStudy (anonymous):

14

OpenStudy (anonymous):

wait no i see

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

@SyedMohammed98 @King.Void.

OpenStudy (anonymous):

5+3+2 = 151022 we observe that 15=5×3 ; 10=5×2 and 22 =15+10−3 9+2+4 = 183652 and 18=9×2 ; 36=9×4; and 52 =18+36−2 8+6+3 = 482466 and 48=8×6; 24=8×3 and 66 =48+24−6 5+4+5 = 202541 and 20=5×4; 25=5×5 and 52 =20+25−4 Therefore for 7+2+5 Now 7×2 =14 ; 7×5=35 and 14+35−2 =47 Hence 7+2+5 = 143547

OpenStudy (anonymous):

they are masters at math so i will let them help you

OpenStudy (brooke2353):

good job @proud_yemeniah_